zarina spent one fourth of her money on food, three tenths of her money on books, and two fifths of her money on records.if she spent $95 on these three items all together, how much money did she have to begin wity?
Answer by mananth(16949)

zarina spent one fourth of her money on food, three tenths of her money on books, and two fifths of her money on records.if she spent $95 on these three
..
let the money she has be $x
x/4 + 3x/10 + 2x/5 = 95
LCD= 20
(5x+6x+8x)/20 = 95
19x= 20*95
19x=1900
x =1900/19
x = $100 the cash she had initially